OpenStudy (sherribelle):

If a quantity of substance decays exponentially, the following equation relates the amount of time t it takes for the substance to lose half its value to the decay rate k: 12Av0=Av0e^kt. a. Solve the equation for t in terms of k. b. What is the half-life of a substance whose decay rate is -7% per year? Round your answer to the nearest tenth of a year. c. Solve the equation for k in terms of t. d. If the half-life of a substance is 7 years, find its decay rate. Round your answer to nearest tenth of a percent.

OpenStudy (sherribelle):

@Isaiah.Feynman

OpenStudy (isaiah.feynman):

\[\frac{ 1 }{ 2 }A_{0} = A_{0}e^{kt}\]

OpenStudy (isaiah.feynman):

Since we are dealing with half lives, that models simply says at some initial time (A_0), a substance has a certain quantity, after it stays past its half life, the new quantity becomes half of the original quantity.

OpenStudy (sherribelle):

In the book, it said it found k before they found t.

OpenStudy (isaiah.feynman):

Okay. I was just explaining the model in your question in words so its easier.

OpenStudy (sherribelle):

Okay, so what's next?

OpenStudy (isaiah.feynman):

We answer the questions. To solve for t in terms of k.

OpenStudy (isaiah.feynman):

We cancel out A to get \[\frac{ 1 }{ 2 } = e^{kt}\]

OpenStudy (sherribelle):

It says there is no solution when I plug in -0.07 for k

OpenStudy (isaiah.feynman):

Okay. We haven't even answered the first question though.

OpenStudy (sherribelle):

I'm getting no solution for that as well.

OpenStudy (isaiah.feynman):

Hang on.

OpenStudy (isaiah.feynman):

So, we use the natural logarithm to bring the t. Like so \[\ln \frac{ 1 }{ 2 } = kt \]

OpenStudy (isaiah.feynman):

So, \[t = \frac{\ln \frac{ 1 }{ 2 } }{ k }\]

OpenStudy (sherribelle):

Then we plug -0.07 in k?

OpenStudy (isaiah.feynman):

We just answered the first question.

OpenStudy (isaiah.feynman):

I found out something that's wrong with the question. If its exponential decay, then the formula should look like this \[\frac{ 1 }{ 2 }A _{0} = A_{0}e^{-kt}\]

OpenStudy (isaiah.feynman):

Which means...\[t = \frac{\ln \frac{ 1 }{ 2 } }{ -k }\]

OpenStudy (isaiah.feynman):

For the second question, the half life if the decay rate is -7% is...\[t = \frac{ \ln\frac{ 1 }{ 2 } }{ -(-0.07) }\] Since k = -0.07